# Position: Interactive Generative Video as Next-Generation Game Engine

### v1

@@ 1 Introduction
During the past year, video generation models have made remarkable progress (OpenAI, 2024; Lab & etc., 2024; Zheng et al., 2024; Chen et al., 2024b; Yang et al., 2024e), demonstrating unprecedented capabilities in large-scale motion dynamics, semantic understanding and concept composition, 3D consistency with physical laws, and long-term temporal coherence in both object structure and appearance. These advances show great potential for effectively simulating real-world physics (OpenAI, 2024; Yang et al., 2024d, 2023), suggesting that these models could serve as capable world models for generating physically plausible videos.
@@
The virtual worlds created by IGV naturally align with video games as they provide interactive environments where players can explore and engage with dynamically generated content, representing a promising direction for next-generation gaming. Recent works (Bruce et al., 2024; DeepMind, 2024; Alonso et al., 2024; Valevski et al., 2024; Che et al., 2024; Feng et al., 2024; Yang et al., 2024b; Yu et al., 2025; Decart, 2024) have demonstrated this potential by training action-conditioned video generation models using action-video pairs collected from classic games like Atari (Alonso et al., 2024), DOOM (Valevski et al., 2024; Yang et al., 2024b), CS:GO (Alonso et al., 2024), Minecraft (Yu et al., 2025; Decart, 2024), and Super Mario Bros (Yang et al., 2024b). These models create interactive gaming experiences by iteratively generating predicted video frames in response to user action inputs.
However, as pointed out by some works (Yu et al., 2025; Feng et al., 2024; DeepMind, 2024), merely replicating existing games through IGV offers limited value over traditional game engines. The revolutionary potential of IGV lies in its ability to create infinite entirely new games through its powerful generative capabilities. Imagine a future where everyone can become a game designer, creating their own games by simply providing design instructions to video generation models, which then generate limitless explorable virtual worlds. This will fundamentally transform both game development and gaming experiences.
In conclusion, this position paper argues that Interactive Generative Video (IGV) serves as the core technology for Generative Game Engine (GGE). GGE will significantly reduce technical barriers in game development while boosting development productivity and creativity through infinite AI-driven content generation.
In this position paper, we first introduce essential preliminary knowledge about video generation and AI-driven game applications in Sec. 2. Sec. 3 analyzes the core capabilities required for next-generation Generative Game Engines (GGE) and demonstrates why Interactive Generative Video (IGV) is uniquely positioned to fulfill these requirements. Sec. 4 presents our comprehensive framework for GGE, providing detailed definitions, analysis, and future prospects for each module within the framework. To guide future research and development, Sec. 5 proposes a hierarchical roadmap that outlines progressive milestones toward fully functional GGE systems. Finally, Sec. 6 and Sec. 7 discuss alternative perspectives and provide concluding remarks.
@@ 2 Preliminaries
2.1 Video Generation Models
Video Generation Models aim to synthesize temporally consistent video sequences $\textbf{x}=\{x_{0},x_{1},\dots,x_{t}\}$, where $x_{i}$ indicates the $i$-th frame. The video is generated by an autoregressive model (Razavi et al., 2019), a diffusion model (Ho et al., 2022), or a masked transformer model (Chang et al., 2022). With the rise of diffusion models (Ho et al., 2020; Song et al., 2021; Song & Ermon, 2019), which have now become the mainstream approach in video generation due to its high-quality generation performance, significant progress has been achieved in this field (Chen et al., 2024b; Blattmann et al., 2023; Zheng et al., 2024; Lab & etc., 2024; Yang et al., 2024e; Peebles & Xie, 2023; OpenAI, 2024).
Conditional Video Generation is formulated as $p(\textbf{x}|c)$ and it varies depending on the type of control signal $c$ which denotes conditions such as text prompts or other control signal. Approaches such as (Guo et al., 2025; Ni et al., 2023; Xing et al., 2023) incorporate images as control signals for the video generator, improving both video quality and temporal relationship modeling. Methods such as Direct-a-Video (Yang et al., 2024c), MotionCtrl (Wang et al., 2024b), and CameraCtrl (He et al., 2024) use camera embedders to adjust camera poses, enabling control over camera movements in generated videos, while (Fu et al., 2025) extends this capability by transforming 2D camera signals into 3D for more advanced control.
Autoregressive Video Generation Models. Since game videos require variable-length or even infinite-length generation to enable interactive game experience, autoregressive mechanism is necessary. An intuitive approach is to adopt GPT-like next-token prediction methods (Kondratyuk et al., 2023; Deng et al., 2024; Wang et al., 2024a); however, this approach often falls short in terms of generation quality. Relying on the exceptional performance of diffusion models, Diffusion Forcing (Chen et al., 2024a) implements autoregression by applying different levels of noise to different frames, allowing the denoising of new frames with higher noise levels while conditioning on previous frames with lower noise levels. Methods (Yu et al., 2025; Feng et al., 2024; Valevski et al., 2024; Decart, 2024) leveraging Diffusion Forcing has achieved remarkable results.
2.2 AI-driven Game Applications
Game Video Generation. Previous works have utilized GANs (Kim et al., 2020, 2021; Menapace et al., 2021) to generate game videos or used NeRF to reconstruct 3D scenes to simulate the game process (Menapace et al., 2022, 2024), but often fall short in terms of generation quality. Using the powerful generative capabilities of diffusion models, some works (Alonso et al., 2024; Valevski et al., 2024; Yang et al., 2024b; Decart, 2024; Feng et al., 2024) have produced high-quality game videos. However, the content is typically confined to specific preexisting games. Open-domain Methods (DeepMind, 2024; Feng et al., 2024; Yu et al., 2025), by utilizing multi-stage training or large-scale training datasets, create new content for video games.
Game Design Assistant. AI-powered design assistants offer numerous advantages throughout the creative process, depending on the tool, the type of AI and the creative workflow. These systems can streamline development, reduce costs, reduce manual effort, improve team collaboration, and even inspire creativity (Liapis, 2015). In the gaming domain, most existing AI-driven design tools primarily assist by auto-completing an ongoing design (Smith et al., 2011) or generating multiple design suggestions for creators to evaluate (Liapis et al., 2014; Migkotzidis & Liapis, 2021; Charity et al., 2022; Torii et al., 2023).
Intelligent Game Agent. Reinforcement learning has long been the predominant approach in this domain. Early efforts explored the use of hierarchical RL (Cai et al., 2023; Jiang et al., 2025; Fan et al., 2022; Lin et al., 2021; Oh et al., 2017; Zhou et al., 2024b; BAAI, 2023) in the context of MineRL competitions (Guss et al., 2019). However, due to the absence of guidance from prior knowledge, such approaches often struggle to perform effectively on long-horizon tasks. With the advancement of large language models (Achiam et al., 2023; Touvron et al., 2023), leveraging their prior knowledge to plan long-horizon tasks has shown promising results. Recent advancements in large language model-related research (Wang et al., 2023a, c; Zhu et al., 2023; Qin et al., 2024b; Yang et al., 2025; Huang et al., 2022) have significantly propelled the progress of agents in long-horizon tasks.
Figure 1: Demonstration of GameFactory (Yu et al., 2025)’s ability to generalize action control capabilities learned from Minecraft data to open-domain scenarios. Examples from its homepage showcase various generalized environments where the learned control mechanisms have been successfully applied.
@@ 3 Why IGV for Generative Game Engine?
Building upon recent advances in video generation, we propose Interactive Generative Video (IGV) as a promising foundation for GGE implementation. IGV extends traditional video generation with three key characteristics that align with GGE requirements: (1) powerful and generalizable generative capabilities for creating unlimited new game content, (2) physics-aware world modeling for realistic interactions, and (3) user-controlled generation for interactive experiences. Beyond these technical alignments, IGV offers another unique advantage: it can leverage the vast amount of existing video data for large-scale model training, enabling robust and diverse content generation.
@@ 3.1 Generalizable Generation for Unlimited Games
Video generation models demonstrate exceptional capabilities in creating not just high-quality visual content, but more importantly, unlimited novel game content that has never existed before. By learning from vast collections of real-world video data, these models develop a deep understanding of visual elements and their relationships, enabling them to generalize beyond specific training game data. Unlike traditional asset creation which is limited to predefined content, these video generation models can continuously generate new, coherent, and diverse game elements by creatively combining and transforming learned patterns. As demonstrated in Figure 1, GameFactory (Yu et al., 2025) is capable of generating action-controllable game videos that can generalize to open-domain scenarios.
@@ 3.2 Physics-aware World Modeling
Figure 2: Physics-aware generation capabilities of video models. Top: Examples from Cosmos (NVIDIA, 2025) demonstrating physical understanding in diverse scenarios including robotics, autonomous driving, manufacturing, and home environments. Bottom: Human motion examples generated by Kling.
Figure 3: Proposed framework of Generative Game Engine (GGE). (a) Framework of GGE shows the architecture and interactions between modules. (b) Technical keywords of each module. The game examples shown in gray boxes demonstrate typical applications of each module’s capabilities, with detailed analysis provided in the main text.
Video generation models demonstrate remarkable potential in understanding the inherent rules of the real world, particularly physical knowledge. During training, to ensure accurate video prediction, these models naturally learn implicit physical priors embedded in training videos. These priors encompass various common physical phenomena, including gravity, elasticity, explosions, collisions, as well as complex motion patterns of humans and animals. While traditional game engines typically rely on predefined physical formulas, extensive manual annotations, or motion capture, IGV leverages its learned physical priors to directly generate physically plausible content. This capability significantly simplifies game engine design and reduces the technical expertise required from developers, thereby enhancing game production efficiency. As shown in Figure 2, the generated video examples demonstrate IGV’s physics-aware capabilities, highlighting its potential value for game development.
@@ 3.3 Interactive Generation with User Control
Current video generation models support various types of control signals that are essential for gaming interactions. These control capabilities are particularly effective in intuitive operations such as camera viewpoint adjustment and character movement control. Such precise and responsive control enables players to naturally interact with the generated content, creating engaging gaming experiences. With rapid development in this field, an increasing number of control signal types are being supported, further expanding interactive possibilities. As shown in Fig. 4, these examples demonstrate IGV’s strong interactive control capabilities and validate its potential for game development.
@@ 3.4 Video Data Accessibility Enables Scaling
Video data offers unique advantages for training generative game engines through its accessibility and unified representation format. Unlike traditional game engines that require various heterogeneous assets (3D models, textures, animations, etc.) with substantial manual effort, videos are widely available across internet platforms and continuously growing through social media and streaming services. Moreover, videos naturally capture diverse real-world phenomena and human experiences, enabling models to learn comprehensive world knowledge through large-scale training. This abundant video data facilitates training powerful video generation models at scale, while using video as a unified representation simplifies the development process by avoiding the complexity of managing different asset formats, making it an ideal foundation for generative game development.
Figure 4: Character control demonstrations from GameNGen (Valevski et al., 2024), showing interactive gameplay operations in generated videos.
@@ 4 Framework of Generative Game Engine
4.1 Overview
We decompose our IGV-centered Generative Game Engine into five functional modules, as illustrated in Figure 3. The framework diagram (a) demonstrates the relationships between these modules, while (b) presents their key technical components and application examples. IGV consists of five core modules: First, the Generation module represents the basic generative capability of the video generation model. Four extension modules are built upon it: the Control module supports different modal control signals and is key to achieving interactivity; the Memory module maintains historical generation content from both dynamic and static aspects, crucial for ensuring temporal consistency; the Dynamics module models the internal rule logic of the game’s virtual world, especially physical rules; while the Intelligence module enables advanced capabilities including causal reasoning and self-evolution. These five modules, through their video interface, create an independent virtual world with its own emergent properties and behaviors. However, a virtual world alone does not constitute a complete game experience, as games require external rules that embody game designers’ intentions, providing players with clear objectives and feedback that create gaming enjoyment. Therefore, we propose an additional GamePlay module based on IGV, which is responsible for implementing these external rule logic within the virtual game world.
@@ 4.2 Generation
❑ Concept. The Generation Module handles video generation, the fundamental functionality of IGV. While ensuring basic video generation requirements like visual quality and motion coherence, this module encompasses three crucial functionalities to achieve optimal interactive experience: (1) Autoregressive Generation enables continuous video synthesis with frame-level control frequency. This supports endless procedural worlds in No Man’s Sky where players can seamlessly explore for hundreds of hours, real-time weather and day-night cycles in Red Dead Redemption 2 that evolve continuously, and instant response to rapid player inputs in rhythm games like Beat Saber where every frame matters. (2) Real-time Processing facilitates low-latency interaction with users. This is essential in competitive games like Counter-Strike, Forza Motorsport, and League of Legends where instant visual feedback is crucial. (3) Multi-modal Generation complements the video output with other modalities like text and audio. This includes dynamic music that responds to gameplay in Journey, positional audio cues for enemy locations in PUBG, ambient sound effects in Minecraft, and real-time dialogue subtitles in Mass Effect.
@@
(1) Autoregressive Generation:
• Diffusion-based methods excel at generating high-quality visual content. A simple way to achieve autoregressive generation is to use different noise levels across frames. Key methods like diffusion forcing (Chen et al., 2024a) and rolling diffusion (Ruhe et al., 2024) have been widely used in game video generation (Feng et al., 2024; Valevski et al., 2024; Decart, 2024; Yu et al., 2025). This approach enables autoregressive generation while preserving visual quality, without any changes to the model structure.
• Next token prediction offers another approach to autoregressive video generation (Kondratyuk et al., 2023; Wang et al., 2024a), though its visual quality currently lags behind diffusion methods. However, its potential for integration with large language models, which could enable strong causal reasoning capabilities (Zhou et al., 2024a; Xie et al., 2024), makes it a promising direction.
• Recent attempts to combine diffusion models with next token prediction aim to maintain high quality while modeling frame-to-frame causality (Li et al., 2024; Deng et al., 2024). While these hybrid approaches show promise, they are still in early stages and their potential to surpass established diffusion-based methods remains to be seen.
@@
From an algorithmic perspective, this can be achieved through various established techniques including distilling to lightweight models (Kim et al., 2025), reducing diffusion sampling steps through ODE distillation (Wang et al., 2023b), using high-compression VAEs or tokenizers (Chen et al., 2024c), and restructuring to causal architectures to avoid redundant computations (Yin et al., 2024). While deployment acceleration strategies are equally important, they are beyond the scope of this discussion.
@@
• One approach is to develop unified large multimodal models that support understanding and generation across multiple modalities including text, vision, audio, human motion, depth maps, and even brain waves. Recent works have started exploring this direction (Zhou et al., 2024a; Xie et al., 2024; Kondratyuk et al., 2023; Reed et al., 2022), though significant challenges remain.
• An alternative strategy is to first develop specialized large models for individual modalities (Deshmukh et al., 2023; Yang et al., 2024a; Jiang et al., 2023; Bai et al., 2024) before integrating them into a unified system. Specifically, this requires designing pipeline relationships between different expert models within the unified system. For example, language models generate video generation instructions, the generated videos then serve as input for audio models to produce corresponding audio signals, ultimately leading to a unified output.
Figure 5: The Control module of IGV manages player control through two aspect: Navigation Control and Interaction Control.

### v2

@@ 1 Introduction
During the past year, video generation models have made remarkable progress sora ; cogvideox ; kling ; veo2 ; seaweed ; moviegen ; runway ; luma ; vidu ; wan ; hunyuanvideo , demonstrating unprecedented capabilities in large-scale motion dynamics, semantic understanding, concept composition, 3D consistency with physical laws, and long-term temporal coherence in both object structure and appearance. These advances show great potential for effectively simulating real-world physics sora ; video-new-language ; unisim , suggesting that these models could serve as capable world models for generating physically plausible videos.
@@
The virtual worlds created by IGV naturally align with video games as they provide interactive environments where players can explore and engage with dynamically generated content, representing a promising direction for next-generation gaming. Recent works genie ; genie2 ; diamond ; gamengine ; gamegenx ; matrix ; playgen ; gamefactory ; oasis ; wham ; adaworld ; mineworld ; maag ; worldmem have demonstrated this potential by training action-conditioned video generation models using action-video pairs collected from classic games like Atari diamond , DOOM gamengine ; playgen , CS:GO diamond , Minecraft gamefactory ; oasis ; mineworld ; worldmem , and Super Mario Bros playgen . These models create interactive gaming experiences by iteratively generating predicted video frames in response to user action inputs.
However, as pointed out by some works gamefactory ; matrix ; genie2 , merely replicating existing games through IGV offers limited value over traditional game engines. The revolutionary potential of IGV lies in its ability to create infinite entirely new games through its powerful generative capabilities. Imagine a future where everyone can become a game designer, creating their own games by simply providing design instructions to video generation models, which then generate limitless explorable virtual worlds. This will fundamentally transform both game development and gaming experiences.
In conclusion, this position paper argues that Interactive Generative Video (IGV) serves as the core technology for Generative Game Engine (GGE). GGE will reduce technical barriers in game development while boosting productivity and creativity through infinite AI-driven content generation.
In this position paper, we first introduce preliminary knowledge about video generation and AI-driven game applications in Sec. 2. Sec. 3 analyzes the core capabilities required for next-generation Generative Game Engines (GGE) and demonstrates why Interactive Generative Video (IGV) is uniquely positioned to fulfill these requirements. Sec. 4 presents our comprehensive framework for GGE, providing detailed definitions, analysis, and future prospects for each module within the framework. To guide future research and development, Sec. 5 proposes a hierarchical roadmap that outlines progressive milestones toward fully functional GGE systems. Finally, Sec.6, Sec.7 and Sec. 8 discuss alternative perspectives, address potential ethical issues and provide concluding remarks.
@@ 2 Preliminaries
A more detailed preliminary section can be found in Appendix A.
Video Generation Models. Video generation models have achieved significant breakthroughs with the rise of diffusion models ddpm ; sde ; score ; flow ; rectified , which have become the mainstream approach due to their superior generation quality sora ; cogvideox ; kling ; veo2 ; seaweed ; moviegen ; runway ; luma ; vidu ; wan ; hunyuanvideo . The field has also made substantial progress in conditional video generation guo2025sparsectrl ; ni2023conditional ; dynamicrafter , particularly in camera control direct-a-video ; motionctrl ; cameractrl ; 3dtrajmaster ; recammaster , where methods like MotionCtrl motionctrl and CameraCtrl cameractrl enable precise manipulation of camera movements. For autoregressive video generation, which is crucial for creating variable-length or infinite video sequences, two representative approaches have emerged: GPT-like next-token prediction methods videopoet ; nova ; emu3 and Diffusion Forcing cdf ; dfot .
AI-driven Game Applications. AI technologies have demonstrated diverse applications in game creation. In game video generation, recent works leveraging diffusion models diamond ; gamengine ; oasis have achieved high-quality results, with open-domain methods genie2 ; matrix ; gamefactory even enabling the creation of novel game content. AI-powered design assistants have enhanced the game development process by automating design completion tanagra and generating multiple design suggestions sketchbook ; susketch , thereby streamlining development and fostering creativity. Furthermore, intelligent game agents have evolved from traditional reinforcement learning approaches gsb ; clip4mc to more sophisticated LLM-based methods voyager ; deps , significantly improving performance in long-horizon tasks.
@@ 3 Why IGV for Generative Game Engine?
Building upon recent advances in video generation, we propose Interactive Generative Video (IGV) as a promising foundation for GGE implementation. As illustrated in Fig. 4 (a), IGV is not the entirety of GGE. From a definitional perspective, IGV is viewed from a technical angle, while GGE is viewed from an application angle. Specifically, IGV represents video generation technology that supports interactive user input control, while GGE represents a game engine that utilizes generative AI to create games. IGV, as a potential realization of GGE, offers four key advantages: (1) powerful generalizable generative capabilities, (2) physics-aware world modeling, (3) user-controlled generation for interactive experiences, and (4) leveraging vast video data for training. In the following subsections, we elaborate on these advantages in detail.
@@ 3.1 Generalizable Generation for Unlimited Games
Figure 1: Demonstration of GameFactory gamefactory ’s ability to generalize action control abilities learned from Minecraft data to open-domain scenarios.
[…]
@@ 3.2 Physics-aware World Modeling
Figure 2: Physics-aware generation capabilities of video models. Top: Examples from Cosmos cosmos demonstrating physical understanding in diverse scenarios including robotics, autonomous driving, manufacturing, and home environments. Bottom: Human motion examples generated by Kling kling .
Video generation models demonstrate remarkable potential in understanding the inherent rules of the real world, particularly physical knowledge sora ; video-new-language ; unisim . During training, to ensure accurate video prediction, these models naturally learn implicit physical priors embedded in training videos. These priors encompass various common physical phenomena, including gravity, elasticity, explosions, collisions, as well as complex motion patterns of humans and animals. While traditional game engines typically rely on predefined physical formulas, extensive manual annotations, or motion capture, IGV leverages its learned physical priors to directly generate physically plausible content. This capability significantly simplifies game engine design and reduces the technical expertise required from developers, thereby enhancing game production efficiency. As shown in Figure 2, the generated video examples demonstrate IGV’s physics-aware capabilities, highlighting its potential value for game development.
Figure 3: GameNGen gamengine shows interactive gameplay in generated videos.
@@ 3.3 Interactive Generation with User Control
Precise control of visual generation models has made significant progress t2iadapter ; controlnet ; control-a-video ; controlnext . Current video generation models support various control signals essential for gaming interactions. These control capabilities excel in intuitive operations such as camera viewpoint adjustment motionctrl ; cameractrl ; 3dtrajmaster ; recammaster and character movement control animate-anyone . Such precise and responsive control enables players to naturally interact with generated content, creating engaging gaming experiences. With rapid development, more control signal types are being supported, further expanding interactive possibilities fulldit ; drag-a-video . Fig. 3 demonstrates IGV’s strong interactive control capabilities and validates its potential for game development.
@@ 3.4 Video Data Accessibility Enables Scaling
Video data offers unique advantages for training generative game engines through its accessibility and unified representation format video-new-language . Unlike traditional game engines that require various heterogeneous assets (3D models, textures, animations, etc.) with substantial manual effort, videos are widely available across internet platforms and continuously growing through social media and streaming services. Moreover, videos naturally capture diverse real-world phenomena and human experiences, enabling models to learn comprehensive world knowledge through large-scale training. This abundant video data facilitates training powerful video generation models at scale, while using video as a unified representation simplifies the development process by avoiding the complexity of managing different asset formats, making it an ideal foundation for generative game development.
@@ 4 Framework of Generative Game Engine
Figure 4: Proposed framework of Generative Game Engine (GGE). (a) Architecture and interactions between modules of GGE. (b) Technical keywords and their explanation of each module.
We decompose our IGV-centered Generative Game Engine into six functional modules. Fig. 4 (a) demonstrates the relationships between these modules, while Fig. 4 (b) presents their key technical components and explanation. IGV consists of five core modules: First, the Generation module represents the basic generative capability of the video generation model. Four extension modules are built upon it: the Control module supports different modal control signals and is key to achieving interactivity; the Memory module maintains historical generation content from both dynamic and static aspects, crucial for ensuring temporal consistency; the Dynamics module models the internal rule logic of the game’s virtual world, especially physical rules; while the Intelligence module enables advanced capabilities including causal reasoning and self-evolution. These five modules, through their video interface, create an independent virtual world with its own emergent properties and behaviors. However, a virtual world alone does not constitute a complete game experience, as games require external rules that embody game designers’ intentions, providing players with clear objectives and feedback that create gaming enjoyment. Therefore, we propose an additional GamePlay module based on IGV, which serves as the key differentiator between GGE and IGV and is responsible for implementing these external rule logic within the virtual game world.
@@ 4.1 Generation
❑ Concept. The Generation Module handles video generation, the fundamental functionality of IGV. While ensuring basic video generation requirements like visual quality and motion coherence, this module encompasses three crucial functionalities to achieve optimal interactive experience: (1) Streaming Generation enables continuous video synthesis with frame-level control frequency. This supports endless procedural worlds in No Man’s Sky where players can seamlessly explore for hundreds of hours, real-time weather and day-night cycles in Red Dead Redemption 2 that evolve continuously, and instant response to rapid player inputs in rhythm games like Beat Saber where every frame matters. (2) Real-time Processing facilitates low-latency interaction with users. This is essential in competitive games like Counter-Strike, Forza Motorsport, and League of Legends where instant visual feedback is crucial. (3) Multi-modal Generation complements the video output with other modalities like text and audio. This includes dynamic music that responds to gameplay in Journey, positional audio cues for enemy locations in PUBG, ambient sound effects in Minecraft, and real-time dialogue subtitles in Mass Effect.
@@
(1) Streaming Generation:
Diffusion-based methods sora excel at generating high-quality visual content. A straightforward way to achieve streaming generation is to use different noise levels across frames. The variable noise levels mechanism means that later frames (with higher noise) can depend on previous frames (with lower noise), implementing autoregressive generation. Representative methods like Diffusion Forcing cdf ; dfot ; causvid ; ar-diffusion ; causalfusion have been widely used in game video generation matrix ; gamengine ; oasis ; gamefactory .
Next token prediction offers another approach to autoregressive video generation videopoet ; emu3 , though its visual quality currently lags behind diffusion methods. However, its potential for integration with LLM, which could enable strong causal reasoning abilities zhou2024transfusion ; xie2024show , makes it a promising direction.
Recent attempts to combine diffusion models with next token prediction aim to maintain quality while modeling frame causality mar ; nova . While these hybrid approaches show promise, they are still in early stages and their potential to surpass established diffusion-based methods remains to be seen.
@@
Figure 5: The Control module manages player control through two aspect: navigation and interaction control.
Recent works have demonstrated promising advances in efficient video generation through various algorithmic techniques. These include lightweight model distillation bk-sdm , ODE-based diffusion step reduction videolcm , high-compression VAEs dcae , and causal architectures like CausVid causvid with distribution matching distillation and Cosmos cosmos with Medusa speculative decoding, key-value caching, and tensor parallelism. These advances, coupled with hardware optimizations like GPU parallelization and quantization, suggest that real-time video generation will soon be accessible to game developers on common hardware. While large companies can provide cloud computing support, we expect future personal developers to run these models on accessible machines.
@@
One approach is to develop unified large multimodal models that support understanding and generation across multiple modalities including text, vision, audio, human motion, depth maps, and so on. Recent works have started exploring this direction zhou2024transfusion ; xie2024show ; videopoet ; gato , though significant challenges remain. An alternative strategy is to first develop specialized large models for individual modalities pengi ; depthanything ; motiongpt ; lvm before integrating them into a unified system. Specifically, this requires designing pipeline relationships between different expert models within the unified system. For example, language models generate video generation instructions, the generated videos then serve as input for audio models to produce corresponding audio signals, ultimately leading to a unified output.